Please help me solve this problem thank you in advance
sol:-since x=2/3 therefore substitute value of x in eq 1 e.g. 3(2/3)+6y=0then 2+6y=0 then 6y=-2 ory=-1/3

Answer by sanjay_jb007(1)

sol:-since x=2/3 therefore substitute value of x in eq 1 e.g. 3(2/3)+6y=0then 2+6y=0 then 6y=-2 ory=-1/3